# Client-wise Modality Selection for Balanced Multi-modal Federated Learning

### v1

@@ 4.2 Modality-level gradient decoupling
$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\nabla\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{P}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{P}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
@@
$+\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{P}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{P}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
@@
Firstly, replacing Eq. 8 with Eq. 9 can enhance the weak modality on the local side. However, the right-hand side of the first equation is still a complex issue. Then, we decouple this objective into two submodular functions according to triangle inequality and the full gradient approximation is divided into two parts: the first part uses selected multi-modal CE gradient to fit fully multi-modal CE gradient aggregation and the second part approximates the fully multi-modal PCE gradient aggregation via selected uni-modal CE gradient and both selected multi- and uni-modal PCE gradient. Two gradient similarity matrices are required for two parts. The modality-level gradient decoupling not only reduces algorithm complexity (convert joint selection to two selection problems) but also highly lowers training cost as each selected client only needs to perform once local training according to its selected modality type to update gradient similarity matrices (both matrices for multi-modal clients, only the second matrix for uni-modal clients).
After modality selection, the gradient of $\theta^{I}$ from each client varies dramatically since there exists an obvious difference between the multi-modal learning gradient and uni-modal learning gradient for the weak modality, as discussed in [9]. To address this heterogeneity, we replace the local prototype in Eq. 10 with the global prototype $c^{GI}$:
$\mathcal{L}_{GP}^{k}\left(v_{I}\right)=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{GI}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{GI}\right)\right)}}\right]$ (12)
Applying the global PCE loss to Eq. 9 can enhance local weak modality and align the feature spaces of various clients simultaneously.
Then, we can perform modality selection with the stochastic greedy algorithm for two submodular functions:
$S_{M}\leftarrow S_{M}\cup k_{1}^{*},$ (13)
$k_{1}^{*}\in$ $\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\right)\right]$
Since the second submodular function involves both $S_{M}$ and $S_{I}$, we add a simple yet effective conflict resolution strategy to ensure $S_{M}\cap S_{I}=\oslash$ as well as, more importantly, balance the learning of different modalities:
$k_{2}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\begin{array}[]{c}\bar{G}\left(S_{M}\cup S_{I}\right)-\\ \bar{G}\left(\left\{k\right\}\cup S_{M}\cup S_{I}\right)\\ \end{array}\right]$
where $\rho_{I}^{k}$ is the imbalanced ratio of client $k$ as in [9], which is the quotient of average ground-truth logits from two modalities (See the Appendix for details). $\rho_{I}^{k}>1$ means audio is dominant while visual is dominant when $\rho_{I}^{k}<1$. $\chi$ is a hyper-parameter for threshold. It tends to select weak modality for severely imbalanced clients to alleviate imbalance, while selected balanced clients use multi-modal data to improve overall performance.
@@
            Aggregate global model, prototypes and $\rho_{I}$;
      Select a subset of modalities $S_{M}$ and $S_{I}$ (or $S_{A}$) using Eqs. 13 and 14;
       Send $\theta$ to selected clients;
@@
             Perform multi-modal learning in $S_{M}$ and uni-modal learning in $S_{I}$ (or $S_{A}$) by Eq. 9 with Eq. 12;
             Send gradients, prototypes and $\rho_{I}^{k}$ to server;
Algorithm 1 CMSFed.
(a) CREMA-D
(b) AVE
(c) CG-MNIST
Figure 3: Test accuracy of CMSFed compared with other baselines on CREMA-D, AVE and CG-MNIST under IID settings. CMSFed converges to more accurate solutions than all baselines in these three datasets.
(a) CREMA-D
(b) AVE
(c) CG-MNIST
Figure 4: Test accuracy of CMSFed compared with other baselines on CREMA-D, AVE and CG-MNIST under non-IID settings. CMSFed still makes remarkable improvement.
Discussion. (1) Mitigating the selection bias in our method are twofold: the PCE loss $\tilde{F}_{k}$ alleviates imbalance at local side and the selected uni-modal clients further promote balanced learning of global model. Meanwhile, the diversity coming from two submodular functions ensures the improvement for the global model. (2) The maximization of $\bar{G}\left(S_{M}\right)$ and $\bar{G}\left(S_{M}\cup S_{I}\right)$ can be viewed as the optimization for multi-modal learning and uni-weak-modal learning respectively. Imbalance ratio is adopted to keep their balance. (3) We assume $I$ is the weak modality above while in practice, we can determine the weak modality before modality selection via the aggregated global imbalance ratio $\rho_{I}$ with $\rho_{I}^{k}$ from all clients. Overall, the pseudo-code of CMSFed is provided in Algorithm 1. (4) Only the gradients, prototypes and $\rho_{I}^{k}$ participate in communication, so there is no privacy issue and similar communication overheads as in traditional FL.
@@ 5.1 Datasets and baselines
CREMA-D [2] is an audio-visual dataset for emotion recognition task, each video in which consists of both facial and acoustic emotional expressions. There are total 6 categories for emotional states: neutral, happy, sad, fear, disgust and anger. 7,442 clips in total are collected in this dataset. 6,698 samples are randomly chosen as the training set and the rest of 744 samples are the testing set.
AVE [30] is an audio-visual dataset for audio-visual event localization, which consists of 28 event classes and 4,143 10-second video clips with both auditory and visual tracks as well as second-level annotations. All the video clips are collected from YouTube. In our experiments, we aim to construct a labeled multi-modal classification dataset by extracting the frames from event-localized video segments and capturing the audio clips within the same segments. The training and validation splits of the dataset follow [30].
Colored-and-gray MNIST [15] is a synthetic dataset based on MNIST [16], and we denote it as CG-MNIST in this paper. Two kinds of images form a sample pair: a gray-scale image and a monochromatic image. In the training set, there are 60,000 sample pairs, and the monochromatic images are strongly color-correlated with their digit labels, In the validation set, the number of sample pairs is 10,000, while the monochromatic images are weakly color-correlated with their labels. The data synthesis method follows [34]. This dataset is used to prove the method’s effectiveness beyond audio-visual modality.
(a) Audio in CREMA-D
(b) Visual in CREMA-D
(c) Audio in AVE
(d) Visual in AVE
Figure 5: The performance of each modality compared with other baselines on CREMA-D and AVE under IID settings. CMSFed achieves comparable audio accuracy and significantly better visual performance simultaneously.
@@ 5.2 Experimental settings
For CREMA-D and AVE, we use ResNet18 [10] as the backbone for both audio and visual modalities. Audio data is converted to a spectrogram of size 257x299 for CREMA-D and 257×1,004 for AVE. We randomly choose 3 frames and 4 frames to build image training sets for CREMA-D and AVE respectively. For CG-MNIST, we build a neural network with 4 convolution layers and 1 average pool layer as the encoder, following the setting as in [9]. We choose the simple yet effective fusion method, concatenation [25], to build fusion classifier for the three datasets. We set 20 clients for CREMA-D, AVE and 30 clients for CG-MNIST, while 5 clients are selected in each communication round for CREMA-D, AVE and 6 for CG-MNIST. For IID setting, training data is uniformly distributed to clients. For non-IID scenarios, we use Dirichlet distribution [11] $Dir\left(\alpha\right)$ to split data ($\alpha=3$ for CREMA-D and AVE, $\alpha=2$ for CG-MNIST). The optimizer is SGD [28] for all datasets. Learning rate is initialized at 1e-3 or 1e-2 for CEAMA-D and AVE or CG-MNIST and becomes 1e-4 or 1e-3 in the later training stage. The hyper-parameter $\chi$ is set to 1.5-2.5 according to datasets and settings. To complete stochastic greedy algorithm for Eqs. 13 and 14, we use the gradients from the selected clients at current round to update part of the similarity matrix, which is named “no-overheads” in [1]. Except for pow-d, DivFL and CMSFed, other baselines select clients randomly. Each client has two modal data by default if not specified.
@@ 5.3 Comparison with baselines
CMSFed effectively improves the performance. As shown in Fig. 3, we report the test accuracy versus the number of communication rounds for three datasets under IID data distributions. We observe that CMSFed can achieve significant improvement compared with all other baselines (by up to 5% on CREMA-D). CMSFed also realizes comparable or even faster convergence speeds in CREMA-D and CG-MNIST. The slow learning at the beginning in AVE is attributed to the need for more effort to explore the information of weak modality. Client sampling here (pow-d and DivFL) cannot fully exploit information for all modalities, making its improvement limited or even worse than FedAvg in CG-MNIST. Although FedOGM and FedPMR accomplish modest improvement because of their ability to alleviate modality imbalance, they are not as good as CMSFed since they do not consider the overall performance of aggregated model. Traditional uni-modal FL methods for statistical heterogeneity (e.g. FedProx) and MFL method for modality incongruity (FedMSplit) only obtain slight improvement. The results for non-IID data distributions are demonstrated in Fig. 4. The overall trends are similar to the results in Fig. 3. These results exhibit the superiority of our method and the importance of meticulously selecting the modalities for training in MFL.
@@
FedAvg-0.2 61.2 58.1 58.5 58.7
FedAvg+ML 55.8 54.5 62.8 60.7
DivFL+ML 57.1 55.6 63.0 61.1
CMSFed-local 63.7 60.3 63.4 60.1
CMSFed 64.5 61.6 64.7 62.1
Table 1: Ablation study. ‘PCE’ denotes using Eq. 9 for local enhancement. ‘CMSFed-local’ uses local prototypes rather than global prototypes.
CMSFed exploits all modalities comprehensively. To show the effect of our method on modality imbalance, we plot the accuracy curves of each modality on CREMA-D and AVE under IID setting. We can see from Fig. 5 that CMSFed could considerably improve the performance of weak modality (visual) and mitigate the modality-level bias as illustrated in Figs. 4(b) and 4(d). Besides, compared with randomly modal abandoning, which significantly reduces audio performance as illustrated in Fig. 1(b), CMSFed achieves comparable audio performance with other baselines. Since audio modality dominates multi-modal learning in these two datasets, pow-d and DivFL selects clients mainly relying on the loss or gradients from audio, which consequently improve audio performance but no improvement on visual modality. FedMSplit mainly focuses on modality incongruity, which is not an issue here, so there is no notable improvement in both modalities. Although FedProto, FedOGM and FedPMR also alleviate the imbalance, FedProto achieves worse audio behavior and FedOGM and FedPMR only focus on local optimization, resulting in the performance gap between them and our CMSFed on the aggregated model, which further indicates that in MFL, it is important to take both local optimization for each modality and the overall performance for aggregated model into consideration simultaneously.

### v2

@@ 3.1 Local Imbalance Alleviation
$F_{k}\left(v_{A},v_{I}\right)=\left\{\begin{array}[]{c}\begin{matrix}\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\gamma^{k}\mathcal{L}_{ME}^{k}\left(v_{A}\right),&\rho_{I}^{k}\leqslant 1\\ \end{matrix}\\ \begin{matrix}\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right),&\rho_{I}^{k}>1\\ \end{matrix}\\ \end{array}\right.$ (8)
where $\rho_{I}^{k}$ indicates the local imbalance ratio for client $k$ calculated based on the modality-wise ground truth prediction, and $\rho_{I}^{k}>1$ means modality $A$ outperforms modality $I$ and vice verse. $\gamma^{k},\beta^{k}\in\left(0,1\right)$ are the modulation coefficients. The calculation details for $\gamma^{k}$, $\beta^{k}$ and $\rho_{I}^{k}$ are in Appendix.
Different from [11], we apply the global prototype instead of local prototypes to aggregate the knowledge from different clients, and the ME loss is not only applied to multi-modal clients, but also to uni-modal clients (See details in Sec. 3.2). Now, the weak modality will be stimulated through local update.
3.2 Balanced Modality Selection
Considering that different types of modal combinations may be selected to participate in local training in our scheme, we define the participated clients as multi-modal clients and uni-modal clients (e.g., client 2 in Fig. 1 Right is a uni-modal client with modal $I$ for training while client 1 and 4 are multi-modal clients with both modalities for training). In Sec. 3.1, we propose the ME loss to facilitate the local learning of weak modality, which is originally designed for multi-modal clients. Here, we also apply it on uni-weak-modal clients to realize gradient consistency on different types of clients.
Assume modality $I$ is weak here. Hence, the local loss for multi-modal and uni-modal clients should be:
$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)$ (9)
$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A}\right)=\mathcal{L}_{CE}^{k}\left(v_{A}\right),F_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)+\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)$
Next, following Eq. 2, we define the paradigm of modality selection for MFL, aiming to approximate the full gradient aggregation from all clients by the gradient from the customized multi-modal and uni-modal clients:
$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}=\sum_{k\in\left[N\right]}{\left[\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\nabla F_{\sigma_{M}\left(k\right)}\left(v_{A},v_{I}\right)\\ -\nabla F_{\sigma_{A}\left(k\right)}\left(v_{A}\right)-\nabla F_{\sigma_{I}\left(k\right)}\left(v_{I}\right)\\ \end{array}\right]}$ (10)
$+\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}+\sum_{k\in S_{A}}{\gamma_{k}^{A}\nabla F_{k}\left(v_{A}\right)}+\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}$
where $\sigma_{M}$, $\sigma_{A}$ and $\sigma_{I}$ map $V\rightarrow S_{M},S_{A},S_{I}$, the client sets who use multi-modal, uni-$A$, and uni-$I$ data for training respectively, and $S_{M}\cap S_{A}=S_{A}\cap S_{I}=S_{M}\cap S_{I}=\oslash$. $\gamma_{k}^{M}$, $\gamma_{k}^{A}$ and $\gamma_{k}^{I}$ have the similar meaning as $\gamma_{k}$ in Eq. 2.
Since modality $I$ is weak here, we omit the uni-$A$ clients as the multi-modal gradient is dominated by modality $A$ [28], which means we do not need to select uni-$A$ clients for its enhancement. Then, bring Eq. 9 to Eq. 10 and follow the operations from Eq. 2 to Eq. 3, we can obtain:
$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\nabla F_{k}\left(v_{A},v_{I}\right)-{\gamma_{i}^{M}}\nabla F_{i}\left(v_{A},v_{I}\right)-{\gamma_{j}^{I}}\nabla F_{j}\left(v_{I}\right)\right\|}$ (11)
$=\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\nabla\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)-\nabla\mathcal{L}_{CE}^{i}\left(v_{A},v_{I}\right)\\ -\nabla\beta^{i}\mathcal{L}_{ME}^{i}\left(v_{I}\right)-\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{ME}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
@@ 3.2 Balanced Modality Selection
$+\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\beta^{k}\mathcal{L}_{ME}^{k}\left(v_{I}\right)-\nabla\beta^{i}\mathcal{L}_{ME}^{i}\left(v_{I}\right)\\ -\nabla\mathcal{L}_{CE}^{j}\left(v_{I}\right)-\nabla\beta^{j}\mathcal{L}_{ME}^{j}\left(v_{I}\right)\\ \end{array}\right\|}$
@@
The right-hand side of the first equation is the modality selection formula that aims to select a group of multi-modal clients $S_{M}$ and uni-modal clients $S_{I}$ to approximate the aggregated gradients from all clients. However, the joint selection for $S_{M}$ and $S_{I}$ is a complex joint optimization problem. Therefore, we decouple this objective into two submodular functions $G\left(S_{M}\right)$ and $G\left(S_{M}\cup S_{I}\right)$ according to triangle inequality, while the full gradient approximation is divided into two parts: the first part uses selected multi-modal CE gradient to fit fully multi-modal CE gradient aggregation, and the second part approximates the fully multi-modal ME gradient aggregation via selected uni-modal CE gradient and both selected multi- and uni-modal ME gradient. The modality-level gradient decoupling converts the complex joint selection to two simply separated selection problems.
@@
            Aggregate global model $\theta$, prototypes $c^{G}$ and $\rho_{I}$;
      Select multi-modality for $S_{M}$ and uni-modality for $S_{I}$ (or $S_{A}$) using Eqs. 12 and 13;
       Send $\theta$, $c^{G}$ and $\rho_{I}$ to selected clients;
@@
             Perform multi-modal learning in $S_{M}$ and uni-modal learning in $S_{I}$ (or $S_{A}$) by with Eqs. 8 and 9;
             Send gradients, local prototypes and $\rho_{I}^{k}$ to server;
Algorithm 1 BMSFed.
Although we can solve the two submodular functions with the stochastic greedy algorithm [24], there are still two issues: (1) the selected client according to $G\left(S_{M}\cup S_{I}\right)$ should be specified whether it is uni-modal client or multi-modal client; (2) the separated selection strategy pays less attention to the global modal bias. To address the two problems, we perform the stochastic greedy algorithm for two submodular functions in parallel and propose a simple yet effective conflict resolution strategy to ensure $S_{M}\cap S_{I}=\oslash$ as well as, more importantly, balance the learning of different modalities on global model:
$S_{M}\leftarrow S_{M}\cup k_{1}^{*},k_{1}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\right)\right]$ (12)
$k_{2}^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S_{M}\backslash S_{I},\mathrm{s}\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S_{M}\cup S_{I}\right)-\bar{G}\left(\left\{k\right\}\cup S_{M}\cup S_{I}\right)\right]$
For every selection, we randomly sample a subset of clients $s$. A multi-modal clients $k_{1}^{*}$ is selected from $s$ according to $G\left(S_{M}\right)$ while $k_{2}^{*}$ is also selected from $s$ according to $G\left(S_{M}\cup S_{I}\right)$. $k_{1}^{*}=k_{2}^{*}$ means using multi-modal data from this client can contribute most to the global model. When $k_{1}^{*}\neq k_{2}^{*}$, we allocate it to uni-modal or multi-modal client according to its local imbalance ratio: if it is severely imbalanced, we use its uni-weak-modal data for training and aggregation to alleviate the global modal bias, otherwise we believe that training with its multi-modal data contributes more than uni-modal data. $\chi$ is a hyper-parameter.
Discussion. (1) Overcoming the modal bias in our method are twofold: the ME loss alleviates imbalance at local side and the selected uni-modal clients further promote balanced learning of global model. Meanwhile, the diversity coming from two submodular functions ensures the representative information for the global model. (2) We assume $I$ is the weak modality above while in practice, we can determine the weak modality before modality selection via the aggregated global imbalance ratio $\rho_{I}=\frac{1}{\sum\nolimits_{k=1}^{N}{n_{k}}}\sum\nolimits_{k=1}^{N}{\rho_{I}^{k}\cdot n_{k}}$. Overall, the pseudo-code of BMSFed is provided in Algorithm 1. (3) Only the gradients, prototypes and $\rho_{I}^{k}$ participate in communication, so there is no privacy issue and similar communication overheads as in traditional FL.
@@ 4.1 Datasets and baselines
Datasets. We conduct experiments on four datasets: (1) CREMA-D [4] is an audio-visual dataset for emotion recognition task with total six categories for emotional states. (2) AVE [33] is an audio-visual dataset for event localization with 28 event classes, and here we use it to construct a labeled multi-modal classification dataset following [11]. (3) Colored-and-gray MNIST (CG-MNIST) [17] is a synthetic dataset based on MNIST [18] with gray-scale and monochromatic images as two modalities, following [38]. (4) ModelNet40 is one of the Princeton ModelNet datasets [39] with 3D objects of 40 categories. The front and back [31] views are considered as two modalities, following [5]. Details of these four datasets are in the Appendix.
@@ 4.2 Experimental settings
For CREMA-D, AVE and ModelNet40, we use ResNet18 [12] as the backbone for audio, visual and flow modalities. Audio data is converted to a spectrogram of size 257x299 for CREMA-D and 257×1,004 for AVE. We randomly choose 3 frames and 4 frames to build image training sets for CREMA-D and AVE respectively. For CG-MNIST, we build a neural network with 4 convolution layers and 1 average pool layer as the encoder, following the setting as in [11]. We choose the simple yet effective fusion method, concatenation [27], to build fusion classifier for all the datasets. We set 20 clients for CREMA-D, AVE and ModelNet40 while the number for CG-MNIST is 30. 5 clients are selected in each communication round for CREMA-D, AVE, ModelNet40 and 6 for CG-MNIST. For IID setting, training data is uniformly distributed to all clients. For non-IID scenarios, we use Dirichlet distribution [13] $Dir\left(\alpha\right)$ to split data ($\alpha=3$ for CREMA-D, AVE, ModelNet40, $\alpha=2$ for CG-MNIST). The optimizer is SGD [30] for all datasets. Learning rate is initialized at 1e-3 or 1e-2 for CEAMA-D, AVE and ModelNet40 or CG-MNIST and becomes 1e-4 or 1e-3 in the later training stage. The hyper-parameter $\chi$ is set to 1.2-2.5 according to datasets and settings. To complete stochastic greedy algorithm for Eqs. 12 and 13, we use the gradients from the selected clients at current round to update part of the similarity matrix, which is named “no-overheads” in [3]. Except for pow-d, DivFL and BMSFed, other baselines select clients randomly. Each client has two modal data by default. We do all experiments on a workstation with an RTX 3090 GPU, a 3.9-GHZ Intel Core i9-12900K CPU and 64GB of RAM.
Table 2: Comparison results on four datasets. The metric is the top-1 accuracy (%). The best is in bold, and the second best is underlined. Our method achieves significant improvement on both IID and non-IID scenarios.
Dataset CREMA-D AVE CG-MNIST ModelNet40
Method IID non-IID IID non-IID IID non-IID IID non-IID
FedAvg 50.7 49.8 62.2 59.7 42.3 41.7 87.2 86.5
FedProx 51.0 49.0 62.6 59.9 42.9 43.6 86.9 87.1
FedProto 58.7 54.0 61.7 58.8 51.5 51.4 87.5 87.2
FedOGM 56.9 56.4 62.8 59.3 57.2 53.0 87.6 87.0
FedPMR 55.5 55.1 63.1 61.6 66.1 63.3 87.6 87.7
pow-d 50.5 50.7 62.5 60.0 41.2 40.3 86.8 86.2
DivFL 51.7 50.8 63.3 59.6 43.0 42.1 86.5 86.4
FedMSplit 52.4 51.6 62.4 60.8 43.5 50.9 87.5 87.4
BMSFed 64.5 61.6 64.7 62.1 70.2 66.7 88.7 87.5
(a) CREMA-D under IID
(b) AVE under IID
(c) CREMA-D under non-IID
(d) AVE under non-IID
Figure 2: Test accuracy of BMSFed compared with other baselines on CREMA-D and AVE. BMSFed converges to more accurate solutions than all baselines.
Table 3: The uni-modal performance comparison on CREMA-D and AVE. The metric is the top-1 accuracy (%). Our method achieves the best uni-visual (the weak modality) performance on all settings and maintains the comparable uni-audio performance.
Dataset CREMA-D AVE
Setting IID non-IID IID non-IID
Method A V A V A V A V
FedAvg 51.2 20.6 50.7 20.2 61.1 26.8 61.4 26.4
FedProx 51.3 20.2 50.1 22.0 60.4 27.1 61.2 26.9
FedProto 50.2 35.3 48.6 39.1 55.7 36.8 59.7 32.8
FedOGM 50.5 35.7 48.8 30.2 58.7 28.8 59.4 29.4
FedPMR 51.5 38.7 50.1 35.9 61.7 39.6 61.7 35.3
pow-d 51.5 20.4 51.6 18.8 61.9 26.9 60.1 27.1
DivFL 52.3 21.1 52.1 22.7 62.7 25.3 61.6 26.3
FedMSplit 52.0 21.8 50.8 21.6 61.3 26.9 62.3 28.7
BMSFed 51.0 41.9 49.3 41.4 59.7 40.2 60.2 38.6
@@ 4.3 Comparison with baselines
BMSFed effectively improves the performance. As demonstrated in Tab. 2, our BMSFed achieves the best results on the four datasets under both IID and non-IID settings (by up to 5.8% on CREMA-D). Client sampling here (pow-d and DivFL) cannot fully exploit information for all modalities, making its improvement limited or even worse than FedAvg in CG-MNIST. Although FedOGM and FedPMR accomplish modest improvement because of their ability to alleviate modality imbalance, they are not as good as BMSFed since they do not consider the overall performance of the global model. Traditional uni-modal FL methods for statistical heterogeneity (e.g. FedProx) and MFL method for modality incongruity (FedMSplit) only obtain slight improvement. We also illustrate the trend of test accuracy versus the number of communication rounds on CREMA-D and AVE in Fig. 2. BMSFed can realize comparable or even faster convergence speeds in CREMA-D and AVE.
BMSFed exploits all modalities comprehensively. To show the effect of our method on addressing modal bias, we report the performance of each modality on CREMA-D and AVE as shown in Tab. 3. The uni-modal performance evaluation follows [11]: a sample is classified into the class corresponding to its nearest prototype. It is clear that BMSFed could considerably improve the performance of weak modality (visual) and mitigate the modality-level bias. Besides, compared with randomly modality abandoning, which significantly reduces audio performance as illustrated in Tab. 1, BMSFed achieves comparable audio performance with other baselines. Although FedProto, FedOGM and FedPMR also alleviate the imbalance, they mainly focus on local optimization, resulting in the performance gap between them and our BMSFed on the aggregated model, which further indicates that in MFL, it is important to take both local optimization for each modality and the overall performance for global model into consideration simultaneously.
Table 4: Ablation study. ‘BMSFed-local’ uses local prototypes rather than global prototypes for ME loss.
@@
FedAvg-0.8 61.2 58.1 58.5 58.7
FedAvg+ME 55.8 54.5 62.8 60.7
DivFL+ME 57.1 55.6 63.0 61.1
BMSFed-local 63.7 60.3 63.4 60.1
BMSFed 64.5 61.6 64.7 62.1